On January 1, 2021, the company obtained a $3 million loan with a 11% interest rate. The building was completed on September 30, 2022. Expenditures on the project were as follows:

January 1, 2021 $ 1,200,000
March 1, 2021 660,000
June 30, 2021 520,000
October 1, 2021 620,000
January 31, 2022 360,000
April 30, 2022 675,000
August 31, 2022 1,080,000


On January 1, 2021, the company obtained a $3 million construction loan with a 11% interest rate. Assume the $3 million loan is not specifically tied to construction of the building. The loan was outstanding all of 2021 and 2022. The company’s other interest-bearing debt included two long-term notes of $4,200,000 and $6,200,000 with interest rates of 6% and 8%, respectively. Both notes were outstanding during all of 2021 and 2022. Interest is paid annually on all debt. The company’s fiscal year-end is December 31.

Required:
1. Calculate the amount of interest that Mason should capitalize in 2021 and 2022 using the weighted-average method.
2. What is the total cost of the building?
3. Calculate the amount of interest expense that will appear in the 2021 and 2022 income statements.
Answer #1
2021 2022
Interest Capitalized $              1,74,066 $   2,40,541
Interest Expense $              9,03,934 $   8,37,459
Total cost of building $            55,29,607
Workings:
Weighted Average rate of all debt:-
$            30,00,000 X 11% = $          3,30,000
$            42,00,000 X 6% = $          2,52,000
$            62,00,000 X 8% = $          4,96,000
$        1,34,00,000 $       10,78,000
Weighted Average rate of all debt = 8.04%
($1078000/ $13400000)
Expenditure for 2021
Jan 1,2021 $            12,00,000 X 12/12 = $       12,00,000
March 1, 2021 $              6,60,000 X 10/12 = $          5,50,000
June 30, 2021 $              5,20,000 X 6/12 = $          2,60,000
October 1, 2021 $              6,20,000 X 3/12 = $          1,55,000
$            30,00,000 $       21,65,000
Interest Capitalized in 2021
$            21,65,000 X 8.04% = $          1,74,066
Expenditure for 2022
Jan 1, 2022 ($3000000 + $174066) $            31,74,066 X 9/9 = $       31,74,066
Jan 31, 2022 $              3,60,000 X 8/9 = $          3,20,000
April 30, 2022 $              6,75,000 X 5/9 = $          3,75,000
August 31. 2022 $            10,80,000 X 1/9 = $          1,20,000
$            52,89,066 $       39,89,066
Interest Capitalized in 2022
$                      39,89,066 X 8.04% X 9/12 = $          2,40,541
Cost of Building
Expenditure for 2021 = $       30,00,000
Interest Capitalized in 2021 = $          1,74,066
Expenditure for 2022 = $       21,15,000
Interest Capitalized in 2022 = $          2,40,541
Total cost of building = $       55,29,607
Interest Expense for 2021:
Total Interest Incurred = $       10,78,000
Less : Interest Capitalized = $          1,74,066
2021 Expense = $          9,03,934
Interest Expense for 2022:
Total Interest Incurred = $       10,78,000
Less : Interest Capitalized = $          2,40,541
2022 Expense = $          8,37,459